OpenStudy (anonymous):

Find all the zeros of the equation -2x^4+27x^2+1200=0

OpenStudy (anonymous):

-3(x^4 - 9x^2 - 400) = 0

OpenStudy (anonymous):

What would the zeros be?

OpenStudy (triciaal):

was that a typo in the original ?

OpenStudy (anonymous):

-3(x^4 - 9x^2 - 400) = 0 -3(x^2 - 25)(x^2 + 16) = 0

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

Yes actually is was a typo oops

OpenStudy (anonymous):

x^2 - 25 = 0 x = 5, -5

OpenStudy (anonymous):

Supposed to be -3 in the beginning

OpenStudy (triciaal):

the highest power of x is four so you will have 4 roots

OpenStudy (anonymous):

x^2 + 16 = 0 x = 4i, -4i

OpenStudy (anonymous):

Because when you set =3 equal to zero: -3 = 0 they are NOT equal to zero!

OpenStudy (anonymous):

But when you simplify, you should write it out as: -3(x^2 + 16)(x-5)(x+5) = 0

OpenStudy (triciaal):

@SavineTwist write out the other 2 as well do not leave the x^2 since they ask for all the zeros

OpenStudy (anonymous):

Thank you